Two loudspeakers emit sound waves along the x-axis. The sound has maximum intensity when the speakers are 22cm apart. The sound intensity decreases as the distance between the speakers is increased, reaching zero at a separation of 50cm .

If the distance between the speakers continues to increase, at what separation will the sound intensity again be a maximum?
Answer #1

Given that,

d = 22 cm (for max intensity)

d' = 50 cm

We need to first determine the wavelength,

\lambda = 2 x (d' - d) = 2 x ( 50 - 22) = 56 cm

As the speaker moves from first to second distance, that is 22 to 50, there occurent a destructive interference(given in ques), so the nest maximum is \lambda/2 after this distance. So

D = d' + \lambda/2 = 50 + 56/2 = 78 cm

Hence, the intensity becomes maximum at D = 78 cm.